# VideoRFT: Incentivizing Video Reasoning Capability in MLLMs via Reinforced Fine-Tuning

### v1

@@ Abstract
Reinforcement fine-tuning (RFT) has shown great promise in achieving human-level reasoning capabilities of Large Language Models (LLMs), and has recently been extended to MLLMs. Nevertheless, reasoning about videos, which is a fundamental aspect of human intelligence, remains a persistent challenge due to the complex logic, temporal and causal structures inherent in video data. To fill this gap, we propose VideoRFT, a novel approach that extends the RFT paradigm to cultivate human-like video reasoning capabilities in MLLMs. VideoRFT follows the standard two-stage scheme in RFT: supervised fine-tuning (SFT) with chain-of-thought (CoT) annotations, followed by reinforcement learning (RL) to improve generalization. A central challenge to achieve this in the video domain lies in the scarcity of large-scale, high-quality video CoT datasets. We address this by building a fully automatic CoT curation pipeline. First, we devise a cognition-inspired prompting strategy to elicit a reasoning LLM to generate preliminary CoTs based solely on rich, structured, and literal representations of video content. Subsequently, these CoTs are revised by a visual-language model conditioned on the actual video, ensuring visual consistency and reducing visual hallucinations. This pipeline results in two new datasets – VideoRFT-CoT-102K for SFT and VideoRFT-RL-310K for RL. To further strength the RL phase, we introduce a novel semantic-consistency reward that explicitly promotes the alignment between textual reasoning with visual evidence. This reward encourages the model to produce coherent, context-aware reasoning outputs grounded in visual input. Extensive experiments show that VideoRFT achieves state-of-the-art performance on six video reasoning benchmarks.
@@ 1 Introduction
The ability to reason about complex videos lies at the core of human cognitive development spelke2007core . Humans, even infants, exhibit a remarkable capacity to understand videos – recognizing what has happened, inferring what will happen next, and explaining why events occur. Replicating this capability in AI systems has become a central goal in video understanding, and has been extensively studied in the field of computer vision over the part decade chen2024mecd ; min2024morevqa ; yi2019clevrer ; TRN . Despite the progress, most AI models remain limited to perceptual-level understanding and struggle to reason about video content with the depth, efficiency, and interpretability characteristic of human cognition.
@@
Recently, the rapid development of video MLLMs, such as Video-ChatGPT VideoChatGPT , VILA VILA , Video-LLaVA VideoLLaVA , have significantly advanced the capabilities of video understanding. However, these models are predominantly answer-driven, i.e., producing answers without explicitly revealing the reasoning process. VoT VoT overcomes this by introducing a human-like reasoning framework that structures the task of video reasoning into five predefined steps using fixed templates. Nevertheless, such a rigid, template-based approach stands in contrast to the flexibility of human cognition, which is able to adapt the reasoning process based on perceptual inputs GroundedCognition ; DualProcessCognition .
@@
In contrast, the very recent advancements, e.g., OpenAI-o1 o1 , DeepSeek-R1 DeepseekR1 , and Kimi-1.5 Kimi15 , have shifted focus towards building LLMs that think before answering. These models show strong proficiency in interpreting complex problems, performing multi-step reasoning, and ultimately arriving at correct answers. A key enabler to such capabilities is reinforcement fine-tuning (RFT) luong2024reft , which typically commences with a warm-up phase of supervised fine-tuning using CoTs, and subsequently refines the model through reinforcement learning algorithms (e.g., PPO PPO , GRPO shao2024deepseekmath ). Beyond the language domain, pioneering efforts have extended RFT to MLLMs to enhance image-based capabilities VisionR1 ; VisualRFT ; ReasonRFT ; R1OneVision ; R1VL ; R1Omni , and some works VideoR1 ; VideoChatR1 ; TinyLLaVA-Video-R1 that concurrently with ours show the potential of RFT in the video domain. However, there is a critical challenge remain unsolved: current video CoT datasets lack the complexity and granularity necessary for advanced video reasoning, which fundamentally limits the ability of models to emulate human-level cognitive capabilities. Moreover, how to ensure that reasoning outputs are faithfully grounded in visual evidence remains underexplored in these works.
@@
Motivated by above analysis, we propose VideoRFT, a novel reinforcement fine-tuning framework to incentive the video reasoning capability in MLLMs (see Fig. 1). To overcome the scarcity of video CoTs, we develop a scalable, cognitively inspired pipeline for high-quality video CoT dataset construction. Specifically, we first extract structured textual descriptions from videos using a MLLM to capture fine-grained visual details. The descriptions are then fed into a reasoning-capable LLM (e.g., DeepSeek-R1), which generates initial CoTs through blind reasoning—relying solely on textual descriptions. However, due to the lack of visual context, these CoTs often suffer from inconsistencies and hallucinations Faithful ; HallucinationSurvey . To resolve this, we develop a cross-modal revision stage, wherein a MLLM refines the initial CoTs by taking original video content, ensuring consistency with visual evidence. Based on this pipeline, we construct two large-scale datasets, i.e., VideoRFT-CoT-102K and VideoRFT-RL-310K, which together support the RFT process in VideoRFT.
@@
Furthermore, to strengthen the RL phase, we develop a novel semantic-consistency reward that explicitly enhances the visual faithfulness of reasoning output in MLLMs. Our key observation is that the reasoning traces of MLLMs are typically structured into three consecutive parts for question parsing, video describing, and abstract reasoning. While the question parsing and abstract reasoning components are not necessarily grounded in the visual input, the video describing part should closely align with the actual visual semantics. Based on this insight, our semantic-consistency reward measures the alignment between the token representations of the video description part and the visual features of the input video. This reward is integrated into the GRPO algorithm to guide MLLMs towards generating visually grounded outputs.
@@ 2.1 Data Collection
We extensively collect video question-answer data covering diverse modalities, task types, and cognitive skills. Given the scarcity of high-quality video data in certain domains (e.g., mathematics, science), we additionally incorporate carefully curated image-based instances. The final dataset contains 310K samples in total, supporting diverse answer formats, including multiple-choice (mc), numerical (num), free-form text generation (free), OCR (ocr), and regression (reg). As shown in Fig. 2, the samples are categorized into five groups according to the type of cognitive skills involving in the reasoning process:
@@ 2.2 Cognitively Inspired CoT Generation
Cognitively Inspired CoT Generation. Given the representation $S_{v}$ and a corresponding question $q$, we invoke a LLM, e.g., DeepSeek-R1 to answer the question and extracts its step-by-step reasoning outputs as the initial CoT, i.e., $\text{CoT}_{v}^{(0)}$:
@@
Here, $P_{\text{cog}}$ denotes a composite prompt consisting of five sub-prompts i.e., $P_{\text{cog}}=[p_{s},p_{t},p_{a},p_{v},p_{r}]$, each guiding a distinct stage of the reasoning trajectory in a manner that mimics human cognitive processing. Concretely, they are defined as follows. (1) Simulated observation prompt ($p_{s}$): Instruct the model to simulate viewing the entire video to form an initial high-level understanding of the video scene. (2) Task understanding prompt ($p_{t}$): Encourage analysis of the question $q$ to infer the task type (e.g., fact, reason, causal relationship). (3) Selective focus prompt ($p_{a}$): Direct attention to specific temporal segments of the video relevant to $q$. (4) Visual reasoning prompt ($p_{v}$): Ground the reasoning process in visual content, encouraging analysis over objects, actions, spatial-temporal relations, and event transitions. (5) Reflective answering prompt ($p_{r}$): Guide the model to derive the final answer, optionally incorporating self-verification or reflection to ensure reasoning quality.
Cross-modal CoT Refinement. A key limitation in the initial CoTs is that it might suffer from visual hallucinations due to the lack of visual cues in Eq. 1. To resolve this issue, we introduce a cross-modal refinement strategy to revise the CoT so that it aligns better with the actual video input. Specifically, we prompt a MLLM, i.e., Qwen2.5-VL Qwen25VL , to compare the initial CoT with the video $v$, identify inconsistencies, and perform necessary revisions:
@@
Finally, we apply a filtering stage to ensure the factual correctness of the resulting CoT annotations. For structured tasks with clear ground-truth labels, we directly discard samples with incorrect final answers. For open-ended tasks, we remove samples exhibiting low semantic consistency (measured by CLIP CLIP ) between the generated answer and the reference answer. This filtering process ensures that the resulting CoT dataset maintains high quality and factual reliability. After filtering, 102K high-confidence samples are retained from the initial pool of 310K, forming VideoRFT-CoT-102K for supervised fine-tuning.
[…]
@@ 2.3 Data Analysis
Fig. 4 presents a comparative analysis of the CoTs in our proposed VideoRFT-CoT-102K and Video-R1 VideoR1 . As shown in Fig. 4 (a), the CoTs in our dataset exhibit a broader and longer token length distribution compared to those in Video-R1 (Fig. 4(c)), indicating that our VideoRFT-CoT-102K contains more elaborate, fine-grained, and nuanced reasoning processes. Additionally, the word cloud in Fig. 4 (b) reveal that CoTs in VideoRFT-CoT-102K are dominated by dynamic, video-centric concepts such as “video”, “main”, “happen”, and “first”. The lexical profile reflects an emphasis on narrative structure and temporal progression, which are the key characteristics of complex video understanding. In contrast, Video-R1 (Fig. 4 (d)) features frequent references to static or declarative content, e.g., “diagram”, “image”, “plant”, and “Earth”, suggesting a stronger bias toward factual descriptions rather than deep reasoning. These results highlight that VideoRFT offers greater expressiveness in reasoning depth, and aligns more closely with the demands of real-world video reasoning tasks. Hence, it provides a better foundation for training video MLLMs with advanced reasoning capabilities.
@@ 3 Video Reinforcement Fine-Tuning
This section presents our approach for video reinforcement fine-tuning as shown in Fig. 5. We first make a brief overview of Group Relative Policy Optimization (GRPO) shao2024deepseekmath in §3.1, and then elaborate on the proposed rule-based reward for efficient reinforcement fine-tuning in §3.2.
@@ 3.1 Group Relative Policy Optimization
GRPO shao2024deepseekmath is a computationally efficient rule-based RL algorithm designed specifically for training large reasoning models. Unlike traditional RL methods such as PPO PPO which require four models (policy, value, reward, and reference), GRPO simplifies the approach by eliminating the value model, significantly reducing memory requirements and training complexity. GRPO operates by generating $K$ multiple candidate responses $\{o_{1},o_{2},\ldots,o_{K}\}$ for each query $q$. These responses are then evaluated using defined reward functions, yielding rewards $\{r_{1},r_{2},\ldots,r_{K}\}$. Afterwards, these rewards are normalized to calculate the advantage $A_{i}$ for each response as:
@@
where $\theta$ denotes model parameters to be updated, $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ are the current and old policy, $\pi_{\text{ref}}$ indicates the reference policy, and $\beta$ is the KL divergence regularization coefficient. The $\epsilon$ is a regularization coefficient, preventing the policy from deviating too far from the reference model.
@@ 3.2.1 Semantic-Consistency Reward
The reward is motivated by the observation that the reasoning trace generated by video MLLMs typically consists of three distinct parts, i.e., question parsing, video describing and abstract reasoning, as shown in Fig. 5. Among them, the video describing stage the model’s understanding of visual content, which is the foundation for subsequent reasoning. Therefore, the reward is designed to selectively promote alignment between this stage and the input video.
Formally, to isolate the video describing sentence from the generated response, we apply a regular expression to locate the first full stop. Empirically, the portion following this punctuation corresponds to the model’s visual interpretation. From this point, we extract a fixed-length span of $M$ tokens $t$, and encode it using SigLIP SIGLIP ’s text encoder: $\bm{t}\!=\!\text{SigLIP}_{\text{text}}(t)$. Additionally, we uniformly sample $F$ frames $\{v^{(0)},\cdots,v^{(F)}\}$ from video $v$, and compute the visual representation of each frame $v^{(i)}$ via SigLIP’s image encoder: $\bm{v}^{(i)}\!=\!\text{SigLIP}_{\text{image}}(v^{(i)})$. Then the final video representation ${\bm{v}}$ is naturally obtained by averaging the frame embeddings: ${\bm{v}}\!=\!\frac{1}{F}\sum_{i=1}^{F}\bm{v}^{(i)}$. We then define the semantic-consistency reward as:
@@
where $\text{cos}(\cdot,\cdot)$ denotes cosine similarity, and $w\!=\!2$ is a scaling constant. The $\max(\cdot,0)$ ensures non-negativity of the reward, while the $\min(\cdot,1)$ stabilizes training by bounding the reward. This stage-aware formulation allows us to reward only the part of reasoning tied to visual comprehension, without penalizing abstract reasoning that appropriately extends beyond the visual scope. The result is enhanced semantic fidelity, reduced hallucinations, and improved alignment during RL.

### v3

@@ Abstract
Reinforcement fine-tuning (RFT) has shown great promise in achieving human-level reasoning capabilities of Large Language Models (LLMs), and has recently been extended to MLLMs. Nevertheless, reasoning about videos, which is a fundamental aspect of human intelligence, remains a persistent challenge due to the complex logic, temporal and causal structures inherent in video data. To fill this gap, we propose VideoRFT, a novel approach that extends the RFT paradigm to cultivate human-like video reasoning capabilities in MLLMs. VideoRFT follows the standard two-stage scheme in RFT: supervised fine-tuning (SFT) with chain-of-thought (CoT) annotations, followed by reinforcement learning (RL) to improve generalization. A central challenge to achieve this in the video domain lies in the scarcity of large-scale, high-quality video CoT datasets. We address this by building a multi-expert, cognition-inspired CoT curation pipeline. First, we devise a cognition-inspired prompting strategy to elicit a reasoning LLM to generate preliminary CoTs based solely on rich, structured, and literal representations of video content. Subsequently, these CoTs are revised by a MLLM conditioned on the actual video, ensuring visual consistency and reducing visual hallucinations. This pipeline results in two new datasets, i.e.VideoRFT-CoT-102K for SFT and VideoRFT-RL-310K for RL. To further strengthen the RL phase, we introduce a novel semantic-consistency reward that explicitly promotes the alignment between textual reasoning and visual evidence. This reward encourages the model to produce coherent, context-aware reasoning outputs grounded in visual input. Extensive experiments show that VideoRFT achieves state-of-the-art performance on six video reasoning benchmarks.
@@ 1 Introduction
The ability to reason about complex videos lies at the core of human cognitive development spelke2007core . Humans, even infants, exhibit a remarkable capacity to understand videos – recognizing what has happened, inferring what will happen next, and explaining why events occur. Replicating this capability in AI systems has become a central goal in video understanding, and has been extensively studied in the field of computer vision over the past decade chen2024mecd ; min2024morevqa ; yi2019clevrer ; TRN . Despite the progress, most AI models remain limited to perceptual-level understanding and struggle to reason about video content with the depth, efficiency, and interpretability that characteristic of human cognition.
@@
Recently, the rapid development of video MLLMs, such as Video-ChatGPT VideoChatGPT , VILA VILA , Video-LLaVA VideoLLaVA , has significantly advanced the capabilities of video understanding. However, these models are predominantly answer-driven, i.e., producing answers without explicitly revealing the reasoning process. VoT VoT overcomes this by introducing a human-like reasoning framework that structures the task of video reasoning into five predefined steps using fixed templates. Nevertheless, such a rigid, template-based approach stands in contrast to the flexibility of human cognition, which is able to adapt the reasoning process based on perceptual inputs GroundedCognition ; DualProcessCognition .
@@
In contrast, the very recent advancements, e.g., OpenAI-o1 o1 , DeepSeek-R1 DeepseekR1 , and Kimi-1.5 Kimi15 , have shifted focus towards building LLMs that think before answering. These models show strong proficiency in interpreting complex problems, performing multi-step reasoning, and ultimately arriving at correct answers. A key enabler to such capabilities is reinforcement fine-tuning (RFT) luong2024reft , which typically commences with a warm-up phase of supervised fine-tuning using CoTs, and subsequently refines the model through reinforcement learning algorithms (e.g., PPO PPO , GRPO shao2024deepseekmath ). Beyond the language domain, pioneering efforts have extended RFT to MLLMs to enhance image-based capabilities VisionR1 ; VisualRFT ; ReasonRFT ; R1OneVision ; R1VL ; R1Omni , and some works VideoR1 ; VideoChatR1 ; TinyLLaVA-Video-R1 tha, concurrently with ours, show the potential of RFT in the video domain. However, there is a critical challenge remaining unsolved: current video CoT datasets lack the complexity and granularity necessary for advanced video reasoning, which fundamentally limits the ability of models to emulate human-level cognitive capabilities. Moreover, how to ensure that reasoning outputs are faithfully grounded in visual evidence remains underexplored in these works.
@@
Motivated by the above analysis, we propose VideoRFT, a novel reinforcement fine-tuning framework to incentivize the video reasoning capability in MLLMs (see Fig. 1). To overcome the scarcity of video CoTs, we develop a scalable, cognitively inspired pipeline that integrates multiple expert models to collaboratively construct high-quality video CoT datasets. Specifically, we first employ a specialized MLLM to extract structured textual descriptions from videos, capturing fine-grained visual details. These descriptions are then processed by a reasoning-capable LLM (e.g., DeepSeek-R1), which generates initial CoTs through blind reasoning—relying solely on textual input. However, due to the lack of direct visual grounding, such CoTs often contain inconsistencies and hallucinations (HallucinationSurvey, ). To mitigate this issue, we introduce a cross-modal revision stage, wherein a MLLM refines the initial CoTs by incorporating the original video, ensuring consistency with visual evidence. Based on this pipeline, we construct two large-scale datasets, i.e., VideoRFT-CoT-102K and VideoRFT-RL-310K, which together support the RFT process in VideoRFT.
@@
Furthermore, to strengthen the RL phase, we develop a novel semantic-consistency reward that explicitly enhances the visual faithfulness of reasoning outputs in MLLMs. Our key observation is that the reasoning traces of MLLMs are typically structured into three consecutive parts for question parsing, video describing, and abstract reasoning. While the question parsing and abstract reasoning components are not necessarily grounded in the visual input, the video describing part should closely align with the actual visual semantics. Based on this insight, our semantic-consistency reward measures the alignment between the token representations of the video description part and the visual features of the input video. This reward is integrated into the GRPO algorithm to guide MLLMs toward generating visually grounded outputs.
@@ 2.1 Data Collection
We extensively collect video question-answer data covering diverse modalities, task types, and cognitive skills. Given the scarcity of high-quality video data in certain domains (e.g., mathematics, science), we additionally incorporate carefully curated image-based instances. The final dataset contains 310K samples in total, supporting diverse answer formats, including multiple-choice (mc), numerical (num), free-form text generation (free), OCR (ocr), and regression (reg). As shown in Fig. 2, the samples are categorized into five groups according to the type of cognitive skills involved in the reasoning process:
@@ 2.2 Cognitively Inspired CoT Generation
Cognitively Inspired CoT Generation. Given the representation $S_{v}$ and a corresponding question $q$, we invoke a LLM, e.g., DeepSeek-R1 to answer the question and extract its step-by-step reasoning outputs as the initial CoT, i.e., $\text{CoT}_{v}^{(0)}$:
@@
Here, $P_{\text{cog}}$ denotes a composite prompt consisting of five sub-prompts i.e., $P_{\text{cog}}=[p_{s},p_{t},p_{a},p_{v},p_{r}]$, each guiding a distinct stage of the reasoning trajectory in a manner that mimics human cognitive processing. Concretely, they are defined as follows. (1) Simulated observation prompt ($p_{s}$): Instruct the model to simulate viewing the entire video to form an initial high-level understanding. (2) Task understanding prompt ($p_{t}$): Encourage analysis of the question $q$ to infer the task type (e.g., fact, reason, causal relationship). (3) Selective focus prompt ($p_{a}$): Direct attention to specific temporal segments of the video relevant to $q$. (4) Visual reasoning prompt ($p_{v}$): Ground the reasoning process in visual content, encouraging analysis over objects, actions, spatial-temporal relations, and event transitions. (5) Reflective answering prompt ($p_{r}$): Guide the model to derive the final answer, optionally incorporating self-verification or reflection to ensure reasoning quality.
Cross-modal CoT Refinement. A key limitation in the initial CoTs is that they might suffer from visual hallucinations due to the lack of visual cues in Eq. 1. To resolve this issue, we introduce a cross-modal refinement strategy to revise the CoT so that it aligns better with the actual video input. Specifically, we prompt a MLLM, i.e., Qwen2.5-VL Qwen25VL , to compare the initial CoT with the video $v$, identify inconsistencies, and perform necessary revisions:
@@
Finally, we apply a filtering stage to ensure the factual correctness of the resulting CoT annotations. For structured tasks with clear ground-truth labels, we directly discard samples with incorrect final answers. For open-ended tasks, we remove samples exhibiting low semantic consistency (measured by CLIP CLIP ) between the generated and the reference answer. This filtering process ensures that the resulting CoT dataset maintains high quality and factual reliability. After filtering, 102K high-confidence samples are retained from the initial pool of 310K, forming VideoRFT-CoT-102K for supervised fine-tuning.
@@ 2.3 Data Analysis
Fig. 4 presents a comparative analysis of the CoTs in our proposed VideoRFT-CoT-102K and Video-R1 VideoR1 . As shown in Fig. 4 (a), the CoTs in our dataset exhibit a broader and longer token length distribution compared to those in Video-R1 (Fig. 4(c)), indicating that our VideoRFT-CoT-102K contains more elaborate, fine-grained, and nuanced reasoning processes. Additionally, the word cloud in Fig. 4 (b) reveals that CoTs in VideoRFT-CoT-102K are dominated by dynamic, video-centric concepts such as “video”, “main”, “happen”, and “first”. The lexical profile reflects an emphasis on narrative structure and temporal progression, which are the key characteristics of complex video understanding. In contrast, Video-R1 (Fig. 4 (d)) features frequent references to static or declarative content, e.g., “diagram”, “image”, “plant”, and “Earth”, suggesting a stronger bias toward factual descriptions rather than deep reasoning. These results highlight that VideoRFT offers greater expressiveness in reasoning depth and aligns more closely with the demands of real-world video reasoning tasks. Hence, it provides a better foundation for training video MLLMs with advanced reasoning capabilities.
@@ 3 Video Reinforcement Fine-Tuning
This section presents our approach for video reinforcement fine-tuning as shown in Fig. 5. We first provide a brief overview of Group Relative Policy Optimization (GRPO) shao2024deepseekmath in §3.1, and then elaborate on the proposed rule-based reward for efficient reinforcement fine-tuning in §3.2.
@@ 3.1 Group Relative Policy Optimization
GRPO shao2024deepseekmath is a computationally efficient rule-based RL algorithm designed specifically for training large reasoning models. Unlike traditional RL methods such as PPO PPO , which require four models (policy, value, reward, and reference), GRPO simplifies the approach by eliminating the value model, significantly reducing memory requirements and training complexity. GRPO operates by generating $K$ candidate responses $\{o_{1},o_{2},\ldots,o_{K}\}$ for each query $q$. These responses are then evaluated using defined reward functions, yielding rewards $\{r_{1},r_{2},\ldots,r_{K}\}$. Afterwards, these rewards are normalized to calculate the advantage $A_{i}$ for each response as:
@@
where $\theta$ denotes model parameters to be updated, $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ are the current and old policy model, $\pi_{\text{ref}}$ indicates the reference policy, and $\beta$ is the KL divergence regularization coefficient. Here, $\epsilon$ is a regularization coefficient, preventing the policy from deviating too far from the reference model.
@@ 3.2.1 Semantic-Consistency Reward
The reward is motivated by the observation that the reasoning trace generated by video MLLMs typically consists of three distinct parts, i.e., question parsing, video describing and abstract reasoning, as shown in Fig. 5. Among them, the video describing stage represents the model’s understanding of visual content, which is the foundation for subsequent reasoning. Therefore, the reward is designed to selectively promote alignment between this stage and the input video.
Formally, to isolate the video describing sentence from the generated response, we apply a regular expression to locate the first full stop. Empirically, the portion following this punctuation corresponds to the model’s visual interpretation. From this point, we extract a fixed-length span of $M$ tokens, denoted $t$, and encode it using SigLIP SIGLIP ’s text encoder: $\bm{t}\!=\!\text{SigLIP}_{\text{text}}(t)$. Additionally, we uniformly sample $F$ frames $\{v^{(0)},\cdots,v^{(F)}\}$ from video $v$, and compute the visual representation of each frame $v^{(i)}$ via SigLIP’s image encoder: $\bm{v}^{(i)}\!=\!\text{SigLIP}_{\text{image}}(v^{(i)})$. Then the final video representation $\bm{v}$} is naturally obtained by averaging the frame embeddings: ${\bm{v}}\!=\!\frac{1}{F}\sum_{i=0}^{F}\bm{v}^{(i)}$. We then define the semantic-consistency reward as:
@@
where $\text{cos}(\cdot,\cdot)$ denotes cosine similarity, and w=2w\!=\!2 is a scaling constant. The $\max(\cdot,0)$ ensures non-negativity of the reward, while the $\min(\cdot,1)$ stabilizes training by bounding the reward. This stage-aware formulation allows us to reward only the part of reasoning tied to visual comprehension, without penalizing abstract reasoning that appropriately extends beyond the visual scope. The result is enhanced semantic fidelity, reduced hallucinations, and improved alignment during RL.